# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

### v1

@@ 3.4 Damping Estimates for $\partial_{r}\phi$
Taking $\partial_{r}$ on (3.8a), we get
$-\rho_{+}\partial_{r}\textnormal{{div}}\psi=\partial_{t}\partial_{r}\phi+u\cdot\nabla\partial_{r}\phi-\partial_{r}f^{0}+[\partial_{r},(\psi\cdot\nabla)]\phi+[\partial_{r},(\tilde{u}\cdot\nabla)]\phi.$
Substituting this into the radial component of equation (3.8b), $\hat{r}\cdot\mathcal{L}(\phi,\psi)=\hat{r}\cdot f$, yields
$\partial_{r}D_{t}\phi+\beta\partial_{r}\phi=g_{1}\qquad\text{ or }\qquad D_{t}\partial_{r}\phi+\beta\partial_{r}\phi=g_{2},$ (3.36)
where $D_{t}\vcentcolon=\partial_{t}+(u\cdot\nabla)$, $\beta\vcentcolon=\rho_{+}^{2}q_{+}/(2\mu+\lambda)>0$, and
$g_{1}\vcentcolon=$ $\partial_{r}f^{0}+\frac{\beta}{q_{+}}\hat{r}\cdot(f-\partial_{t}\psi)+\frac{\mu\beta}{\rho_{+}q_{+}}\big{\{}\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi\big{\}},$
$g_{2}\vcentcolon=$ $g_{1}+[(\psi\cdot\nabla),\partial_{r}]\phi+[(\tilde{u}\cdot\nabla),\partial_{r}]\phi.$
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $T\in[0,T_{0}]$, and $k+l=0,1,2$,
$\sup\limits_{0\leq t\leq T}\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}\phi\big{\|}^{2}+|u_{b}|\iint_{\tilde{\Gamma}_{T}}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$+\int_{0}^{T}\!\Big{\{}\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}\phi\big{\|}^{2}+\big{\|}\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}D_{t}\phi\big{\|}^{2}\Big{\}}\,{\mathrm{d}}t$
$\leq$ $C\big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big{\}}+C\!\!\int_{0}^{T}\!\!\!\Big{\{}\sum_{j=0}^{k+1}\big{\|}\sqrt{\chi}D\partial^{j}\psi\big{\|}_{l}^{2}+\big{\|}\sqrt{\chi}\partial_{r}^{l}\partial^{k}\partial_{t}\psi\big{\|}^{2}\Big{\}}\,{\mathrm{d}}t.$
@@
Taking derivative $\partial_{r}^{l}\partial^{k}$ on the first equation in (3.36), multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}D_{t}\phi$, and using continuity equation (1.1a), we get
@@
$=$ $\beta\chi\rho\partial_{r}^{l+1}\partial^{k}\phi[D_{t},\partial_{r}^{l+1}\partial^{k}]\phi+\chi\rho\partial_{r}^{l}\partial^{k}g_{1}\partial_{r}^{l+1}\partial^{k}D_{t}\phi+\beta\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
In addition, taking derivative $\partial_{r}^{l}\partial^{k}$ on the second equation in (3.36), multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi$, and using continuity equation (1.1a), we get
@@
$=$ $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi\big{\{}\partial_{r}^{l}\partial^{k}g_{2}+[D_{t}\partial_{r}\,,\partial_{r}^{l}\partial^{k}]\phi\big{\}}+\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
Adding the above two equalities together, integrating the resultant equality in $(t,x)\in[0,T]\times\Omega$, then it follows from boundary condition (1.2) and Cauchy-Schwarz’s inequality that
@@
$+C\iint_{Q_{T}}\chi\Big{\{}\big{|}\partial_{r}^{l}\partial^{k}[(\psi\cdot\nabla),\partial_{r}]\phi\big{|}^{2}+\big{|}\partial_{r}^{l}\partial^{k}[(\tilde{u}\cdot\nabla),\partial_{r}]\phi\big{|}^{2}\Big{\}}\,{\mathrm{d}}x{\mathrm{d}}t$
@@
Using the inequality (3.12), Proposition B.1, and the fact that $\partial_{t}$, $\partial_{r}$, $\partial$ commute with each other, we obtain
@@
By the definition (3.8c)–(3.8d), and inequality (3.12), we obtain
@@
By Proposition B.2, it follows that
@@
Using Lemma 1.1, and inequality (3.12), we get
@@
Putting the estimates for $\mathcal{J}_{1}$$\mathcal{J}_{5}$ into (3.37), we obtain the desired estimate. $\square$
@@ 3.5 Estimates from elliptic and Stokes systems
Let $\nu_{1},\,\nu_{2}>0$. Given a vector-valued function :$\mathcal{R}\vcentcolon\Omega\to\mathbb{R}^{3}$, which belongs to some Sobolev space. Suppose $\zeta\vcentcolon\Omega\to\mathbb{R}^{3}$ is a solution to the elliptic system:
$-\nu_{1}\Delta\zeta^{i}-\nu_{2}\partial_{x_{i}}\textnormal{{div}}\zeta=$ $\mathcal{R}^{i}$ for $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (3.38a)
$\zeta(x)=$ 00 0 for $\text{for }\ x\in\partial\Omega.$ (3.38b)
@@
In addition, suppose $\alpha\vcentcolon\partial\Omega\to\mathbb{R}^{3}$, $h^{0}\vcentcolon\Omega\to\mathbb{R}$, and $h\vcentcolon\Omega\to\mathbb{R}^{3}$ are functions which belong to some Sobolev spaces. Let $w\vcentcolon\Omega\to\mathbb{R}$ and $v\vcentcolon\Omega\to\mathbb{R}^{3}$ be solution to the Stokes-type equations:
$\rho_{+}\textnormal{{div}}v$ $=h^{0}$ for $\text{for }\ x\in\Omega,$ (3.39a)
$-\mu\Delta v^{i}+\rho_{+}q_{+}\partial_{x_{i}}w$ $=h^{i}$ for $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (3.39b)
$v(x)$ $=\alpha(x)$ for $\text{for }\ x\in\partial\Omega.$ (3.39c)
@@
The proof for the above lemma can be found in [15] (See Lemma 4.3). We also note that for the incompressible case: $h^{0}\equiv 0$, the system (3.39) reduces to the well-known Stokes equations, and the similar estimates can be found in [2, 5].
@@
Let $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$ and $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$. For $l=0,1,2$,
@@
For $k\geq 1$ and $k+l=1,2$, the following estimate holds
@@
Proof.
Applying the operator $\chi\partial^{k}$ on the equations (3.8a)–(3.8b), we have that $(v,w)\equiv(\chi\partial^{k}\phi,\chi\partial^{k}\psi)$ solves the Stokes-type system (3.39) with boundary condition $\chi\partial^{k}\psi|_{\partial\Omega}=0\equiv\alpha$, and
By (3.30), we have $|(\nabla\chi\cdot\nabla)F|\leq C\sqrt{\chi}|\partial_{\theta}F|$. Then by Lemma 3.11 and Propositions B.1, we have
$+C\Big{\{}\sum_{m=1}^{k}\big{\|}\partial^{m}\psi\big{\|}_{H^{l+1}(\tilde{\Omega})}^{2}+\sum_{m=0}^{k-1}\big{\|}D\partial^{m}\phi\big{\|}_{H^{l}(\tilde{\Omega})}^{2}\Big{\}}$
This proves the second inequality. The first statement is shown by the exact same argument by setting $v=\psi$, $w=\phi$ in Lemma 3.10. $\square$
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for all $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
@@
Throughout this proof, we use the abbreviation:
$M_{\delta}(T)\vcentcolon=\frac{C}{\delta^{7}}\Big{\{}N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big{\}}+C\delta N^{2}(T).$ (3.40)
First, set $(k,l)=(0,0)$ and $(k,l)=(1,0)$ in Lemma 3.8. Then using Lemmas 3.4, 3.5, 3.6, and Corollary 3.7, we get
@@
Set $(k,l)=(0,0)$ in Lemma 3.11. Then by Lemma 3.5, Corollary 3.7, and (3.41),
@@
Set $(k,l)=(1,0)$ in Lemma 3.11. Then by Lemmas 3.4, 3.5, 3.6, and (3.41)–(3.42),
@@
Set $(k,l)=(0,1)$ in Lemma 3.8. Then using Lemmas 3.4, 3.5, 3.6, Corollary 3.7, and (3.42)–(3.43), we obtain
@@
Set $(k,l)=(0,1)$ in Lemma 3.11. Then using Lemmas 3.4, 3.5, Corollary 3.7, inequalities (3.41) and (3.44), we have
@@
Next, according to equation (3.8b), we set $\zeta\equiv\psi$ in Lemma 3.9. Then by Lemma 3.5, and (3.41), it follows that
@@
Finally, taking derivative $D$ on the equation (3.8a), we get $\partial_{t}D\phi=Df^{0}-D(u\cdot\nabla\phi)-\rho_{+}D\textnormal{{div}}\psi$. Using (3.12) and (3.42), we get that $\|\partial_{t}\phi\|_{L^{2}(0,T;H^{1}(\Omega))}\leq M_{\delta}(T)$. $\square$
@@
There exists a constant $C\!=\!C(\rho_{+},\gamma,K,\mu,\lambda)>\!0$ such that for $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
@@
For simplicity, we use the same notation $M_{\delta}(T)$ as defined in (3.40). Taking $\partial_{t}$ on the equation (3.8b), we verify that $\zeta\equiv\partial_{t}\psi$ solves the elliptic system (3.38), with $\nu_{1}\equiv\mu$, $\nu_{2}\equiv\mu+\lambda$ and
$\mathcal{R}^{i}\equiv\rho_{+}\partial_{t}f^{i}-\rho_{+}\partial_{t}^{2}\psi^{i}-\rho_{+}(\partial_{t}\psi\cdot\nabla)\psi^{i}-\rho_{+}(\psi\cdot\nabla)\partial_{t}\psi^{i}-\rho_{+}(\tilde{u}\cdot\nabla)\partial_{t}\psi^{i}-q_{+}\partial_{x_{i}}\partial_{t}\phi.$
Thus applying Lemma 3.9 on $\zeta\equiv\partial_{t}\psi$ with $k=0$, then using inequality (3.12), and Lemmas 3.5, 3.12, we obtain that
@@
Take $(k,l)=(2,0)$ in Lemma 3.8. Then by Lemmas 3.4, 3.6, Corollary 3.7, and (3.45),
@@
Take $(k,l)=(2,0)$ in Lemma 3.11. Then by Lemmas 3.6, 3.12 and (3.45)–(3.46),
@@
Take $(k,l)=(1,1)$ in Lemma 3.8. Then by Lemmas 3.6, 3.12, Corollary 3.7, inequalities (3.45), and (3.47), we have
@@
Take $(k,l)=(0,2)$ in Lemma 3.8. Then by Lemmas 3.6, 3.12, Corollary 3.7, inequalities (3.45) and (3.49), we obtain
@@
Finally, from the momentum equation (3.8b), we see that $\zeta\equiv\psi$ solves system (3.38) with $\nu_{1}\equiv\mu$, $\nu_{2}\equiv\mu+\lambda$, and $\mathcal{R}\equiv\rho_{+}\{f-\partial_{t}\psi-(u\cdot\nabla)\psi-q_{+}\nabla\phi\}$. Applying Lemma 3.9 on $\zeta\equiv\psi$ with $k=3$, we get $\sup_{0\leq t\leq T}\|D^{3}\psi(t,\cdot)\|^{2}\leq M_{\delta}(T)$. $\square$
@@
By (3.9) and Lemmas 3.123.13, there exists a positive constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)$ such that for all $\delta\in(0,1]$ and $T\in[0,T_{0}]$,
@@
$N^{2}(T)\leq 4^{7}C^{8}N^{2}(0)+\frac{3}{4}N^{2}(T)\qquad\text{ for all }\ T\in[0,T_{0}].$
Therefore $N^{2}(T_{0})\leq 4^{8}C^{8}N^{2}(0)$, and we conclude the proof of Theorem 3.2. $\square$
Appendix A Hardy’s Inequality

### v2

@@ 2.5 Damping estimates for $\partial_{r}\phi$
In this subsection, we derive estimates involving the radial derivative of $\phi$. This is obtained by exploiting the coupling structure of the continuity equation (2.1b) and momentum equation (2.1c). More specifically, we show that $\partial_{r}\phi$ solves a first-order hyperbolic equation via the following procedure: first, we take $\partial_{r}$ on (2.1b) to get
$-\rho_{+}\partial_{r}\textnormal{{div}}\psi=$ $\partial_{r}D_{t}\phi-\partial_{r}f^{0},$ (2.30)
where $D_{t}\vcentcolon=\partial_{t}+(u\cdot\nabla)$. Next, taking the radial component of (2.1c), we obtain
$\frac{2\mu+\lambda}{\rho_{+}}\partial_{r}\textnormal{{div}}\psi+q_{+}\partial_{r}\phi=\hat{r}\cdot(f-\partial_{t}\psi)+\frac{\mu}{\rho_{+}}\big\{\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi\big\}.$
Substituting (2.30) in the above equation yields
$\partial_{r}D_{t}\phi+\beta\partial_{r}\phi=g,$ (2.31)
where
$\beta\vcentcolon=\frac{\rho_{+}^{2}q_{+}}{2\mu+\lambda}>0,\qquad g\vcentcolon=\partial_{r}f^{0}+\frac{\beta}{q_{+}}\hat{r}\cdot(f-\partial_{t}\psi)+\frac{\mu\beta}{\rho_{+}q_{+}}\big\{\hat{r}\cdot\Delta\psi-\partial_{r}\textnormal{{div}}\psi\big\}.$
Furthermore, computing the commutator $\partial_{r},D_{t}]$, we also obtain that
$D_{t}\partial_{r}\phi+\beta\partial_{r}\phi=g+\partial_{r}u\cdot\nabla\phi.$ (2.32)
The term $\beta\partial_{r}\phi$ on the left hand side indicates the dissipative damping effect on $\partial_{r}\phi$.
@@
Suppose $(\phi,\psi)\in X(T_{0})$ is a solution to the system (2.1), where $T_{0}>0$ is the time given in (2.3)–(2.4). Then there exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for an arbitrary $T\in[0,T_{0}]$, it holds that
$\sup\limits_{0\leq t\leq T}\big\|\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}\phi\big\|^{2}+\int_{0}^{T}\!\!\big\|\sqrt{\chi}\partial_{r}^{l+1}\partial^{k}(\phi,D_{t}\phi)\big\|^{2}\,{\mathrm{d}}t+|u_{b}|\iint_{\tilde{\Gamma}_{T}}\!\!|\partial_{r}^{l+1}\partial^{k}\phi|^{2}\,{\mathrm{d}}x{\mathrm{d}}t$
$\leq$ $C\big\{N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\big\}+C\!\!\int_{0}^{T}\!\!\!\Big\{\sum_{j=0}^{k+1}\big\|\sqrt{\chi}D\partial^{j}\psi\big\|_{l}^{2}+\big\|\sqrt{\chi}\partial_{r}^{l}\partial^{k}\partial_{t}\psi\big\|^{2}\Big\}\,{\mathrm{d}}t,$
where $k$, $l$ are non-negative integers such that $k+l=0,1,2$.
@@
Taking the derivative $\partial_{r}^{l}\partial^{k}$ on (2.31) for $k+l=0,1,2$, multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}D_{t}\phi$, and using continuity equation (1.1a), we get
@@
$=$ $\beta\chi\rho\partial_{r}^{l+1}\partial^{k}\phi[D_{t},\partial_{r}^{l+1}\partial^{k}]\phi+\chi\rho\partial_{r}^{l}\partial^{k}g\partial_{r}^{l+1}\partial^{k}D_{t}\phi+\beta\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
In addition, taking the derivative $\partial_{r}^{l}\partial^{k}$ on (2.32), multiplying both sides by $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi$, and using continuity equation (1.1a), we get
@@
$=$ $\chi\rho\partial_{r}^{l+1}\partial^{k}\phi\Big\{\partial_{r}^{l}\partial^{k}g+\partial_{r}^{l}\partial^{k}\big(\partial_{r}u\cdot\nabla\phi\big)+[D_{t}\partial_{r}\,,\partial_{r}^{l}\partial^{k}]\phi\Big\}+\frac{\rho}{2}|\partial_{r}^{l+1}\partial^{k}\phi|^{2}(u\cdot\nabla)\chi.$
Summing the above equations, integrating the resultant equality in $(t,x)\in[0,T]\times\Omega$, it follows from boundary condition (1.2) and Cauchy-Schwarz’s inequality that
@@
$+C\iint_{Q_{T}}\chi\Big\{\big|\partial_{r}^{l}\partial^{k}(\partial_{r}\psi\cdot\nabla\phi)\big|^{2}+\big|\partial_{r}^{l}\partial^{k}(\partial_{r}\tilde{u}\cdot\nabla\phi)\big|^{2}\Big\}\,{\mathrm{d}}x{\mathrm{d}}t$
@@
Using the inequality (2.5), Proposition A.2, and the fact that $\partial_{t}$, $\partial_{r}$, $\partial$ commute with each other, we obtain
@@
By the definition (2.1d)–(2.1e) and inequality (2.5), we obtain
@@
Owing to Proposition A.3, it follows that
@@
Using Lemma 1.1 and inequality (2.5), we get
@@
Substituting the estimates for $\mathcal{J}_{1}$$\mathcal{J}_{5}$ in (2.33), we obtain the desired estimate. $\square$
@@ 2.6 Estimates from elliptic and Stokes systems
Using the previous results derived in Sections 2.32.5, we obtain the estimates containing full spatial derivatives $\|(\nabla^{k}\psi,\nabla^{k}\phi)\|$ for $k=1,2,3$. This is achieved by reformulating equations (2.1b)–(2.1c) and applying elliptic and Stokes-type estimates. To illustrate this, we first present the two known results for solutions to the elliptic system and Stokes’ equations.
Let $\nu_{1},\,\nu_{2}>0$. Given a vector-valued function $\mathcal{R}\vcentcolon\Omega\to\mathbb{R}^{3}$, which belongs to $H^{m}(\Omega)$ for $m=0,1$. Suppose $\zeta\vcentcolon\Omega\to\mathbb{R}^{3}$ is a solution to the elliptic system:
$-\nu_{1}\Delta\zeta^{i}-\nu_{2}\partial_{x_{i}}\textnormal{{div}}\zeta=$ $\mathcal{R}^{i}$ $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (2.34a)
$\zeta(x)=$ $0$ $\text{for }\ x\in\partial\Omega.$ (2.34b)
@@
In addition, suppose $\alpha\vcentcolon\partial\Omega\to\mathbb{R}^{3}$, $h^{0}\vcentcolon\Omega\to\mathbb{R}$, and $h\vcentcolon\Omega\to\mathbb{R}^{3}$ are functions which belong to the suitable Sobolev spaces. Let $w\vcentcolon\Omega\to\mathbb{R}$ and $v\vcentcolon\Omega\to\mathbb{R}^{3}$ be solution to the Stokes-type equations:
$\rho_{+}\textnormal{{div}}v$ $=h^{0}$ $\text{for }\ x\in\Omega,$ (2.35a)
$-\mu\Delta v^{i}+\rho_{+}q_{+}\partial_{x_{i}}w$ $=h^{i}$ $\text{for }\ x\in\Omega\ \text{ and }\ i=1,2,3,$ (2.35b)
$v(x)$ $=\alpha(x)$ $\text{for }\ x\in\partial\Omega.$ (2.35c)
@@
The proof for the above lemma is presented in [19] (See Lemma 4.3). We also note that for the incompressible case: $h^{0}\vcentcolon=0$, the system (2.35) reduces to the well-known Stokes equations, and the similar estimates are derived in [4, 8].
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $l=0,1,2$,
@@
Moreover, for $k\geq 1$ and $k+l=1,2$,
$\big\|\chi D^{2+l}\partial^{k}\psi\big\|^{2}+\big\|\chi D^{1+l}\partial^{k}\phi\big\|^{2}\!\leq$ $C\Big\{\big\|\chi\partial^{k}D_{t}\phi\big\|_{l+1}^{2}\!\!+\big\|\chi\partial_{t}\partial^{k}\psi\big\|_{l}^{2}\!+\big\|f^{0}\big\|_{k+l+1}^{2}\!\!+\big\|f\big\|_{k+l}^{2}\Big\}$
$+C\Big\{\sum_{m=1}^{k}\big\|\partial^{m}\psi\big\|_{H^{l+1}(\tilde{\Omega})}^{2}+\sum_{m=0}^{k-1}\big\|D\partial^{m}\phi\big\|_{H^{l}(\tilde{\Omega})}^{2}\Big\},$
where $\tilde{\Omega}$ is the domain defined in (2.24).
Proof.
Applying the operator $\chi\partial^{k}$ on the equations (2.1b)–(2.1c), we have that $(v,w)=(\chi\partial^{k}\phi,\chi\partial^{k}\psi)$ solves the Stokes-type system (2.35) with $\chi\partial^{k}\psi|_{\partial\Omega}=0=\alpha$, and
By (2.23), we have $|(\nabla\chi\cdot\nabla)F|\leq C\sqrt{\chi}|\partial_{\theta}F|$. Then by Lemma 2.11 and Propositions A.2, we have
@@
This proves the second inequality. The first statement is shown by the exact same argument except we set $v=\psi$, $w=\phi$ in Lemma 2.10. $\square$
@@
There exists a constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ such that for $\delta\in(0,1)$ and $T\in[0,T_{0}]$,
@@
Throughout this proof, we use the abbreviation
$M_{\delta}(T)\vcentcolon=\frac{C}{\delta^{7}}\Big\{N^{2}(0)+|u_{b}|N^{2}(T)+N^{3}(T)\Big\}+C\delta N^{2}(T),$ (2.36)
where $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ is a constant which is independent of $T\in[0,T_{0}]$, $|u_{b}|$, $\delta\in(0,1)$ and the initial data $(\rho_{0},u_{0})$.
First, set $(k,l)=(0,0)$ and $(k,l)=(1,0)$ in Lemma 2.8. Then using Lemmas 2.4, 2.5, 2.6, and Corollary 2.7, we get
@@
Let $(k,l)=(0,0)$ in Lemma 2.11. Then by Lemma 2.5, Corollary 2.7, and (2.37),
@@
Let $(k,l)=(1,0)$ in Lemma 2.11. Then by Lemmas 2.4, 2.5, 2.6, and (2.37)–(2.38),
@@
Let $(k,l)=(0,1)$ in Lemma 2.8. Then using Lemmas 2.4, 2.5, 2.6, Corollary 2.7, and (2.38)–(2.39), we obtain
@@
Let $(k,l)=(0,1)$ in Lemma 2.11. Then using Lemmas 2.4, 2.5, Corollary 2.7, inequalities (2.37) and (2.40), we have
@@
Next, according to equation (2.1c), we set $\zeta=\psi$ in Lemma 2.9. Then by Lemma 2.5, and (2.37), it follows that
@@
Since $D^{2}\partial_{t}\phi=D^{2}D_{t}\phi-D^{2}u\cdot\nabla\phi-2Du\cdot\nabla\phi-u\cdot D^{2}\nabla\phi$, (2.5) and (2.40) implies that
$\int_{0}^{T}\|\partial_{t}\phi\|_{2}^{2}\,{\mathrm{d}}t\leq M_{\delta}(T).$
This completes the proof. $\square$
@@
There exists a constant $C\!=\!C(\rho_{+},\gamma,K,\mu,\lambda)>\!0$ such that for $\delta\in(0,1)$ and $T\in[0,T_{0}]$,
@@
For simplicity, we use the same notation $M_{\delta}(T)$ as defined in (2.36). Taking the temporal derivative $\partial_{t}$ on the equation (2.1c), we verify that $\zeta=\partial_{t}\psi$ solves the elliptic system (2.34), with $\nu_{1}=\mu$, $\nu_{2}=\mu+\lambda$ and
$\mathcal{R}^{i}=\rho_{+}\partial_{t}f^{i}-\rho_{+}\partial_{t}^{2}\psi^{i}-\rho_{+}(\partial_{t}\psi\cdot\nabla)\psi^{i}-\rho_{+}(\psi\cdot\nabla)\partial_{t}\psi^{i}-\rho_{+}(\tilde{u}\cdot\nabla)\partial_{t}\psi^{i}-q_{+}\partial_{x_{i}}\partial_{t}\phi.$
Thus applying Lemma 2.9 on $\zeta=\partial_{t}\psi$ with $k=0$, and using inequality (2.5), and Lemmas 2.5, 2.12, we obtain that
@@
Take $(k,l)=(2,0)$ in Lemma 2.8. Then by Lemmas 2.4, 2.6, Corollary 2.7, and (2.41),
@@
Take $(k,l)=(2,0)$ in Lemma 2.11. Then by Lemmas 2.6, 2.12 and (2.41)–(2.42),
@@
Take $(k,l)=(1,1)$ in Lemma 2.8. Then by Lemmas 2.6, 2.12, Corollary 2.7, inequalities (2.41), and (2.43), we have
@@
Take $(k,l)=(0,2)$ in Lemma 2.8. Then by Lemmas 2.6, 2.12, Corollary 2.7, inequalities (2.41) and (2.45), we obtain
@@
Finally, from the momentum equation (2.1c), we see that $\zeta=\psi$ solves system (2.34) with $\nu_{1}=\mu$, $\nu_{2}=\mu+\lambda$, and $\mathcal{R}=\rho_{+}\{f-\partial_{t}\psi-(u\cdot\nabla)\psi-q_{+}\nabla\phi\}$. Applying Lemma 2.9 on $\zeta=\psi$ with $k=3$, we get $\sup_{0\leq t\leq T}\|D^{3}\psi(t,\cdot)\|^{2}\leq M_{\delta}(T)$. $\square$
@@
By (2.2) and Lemmas 2.122.13, there exists a positive constant $C=C(\rho_{+},\gamma,K,\mu,\lambda)$ such that for $\delta\in(0,1)$ and $t\in[0,T]$,
@@
$N^{2}(t)\leq 4^{7}C^{8}N^{2}(0)+\frac{3}{4}N^{2}(t)\qquad\text{ for all }\ t\in[0,T].$
Therefore $N^{2}(T)\leq 4^{8}C^{8}N^{2}(0)$, and we conclude the proof of Theorem 2.2. $\square$
Declaration of competing interests
The authors declare that there is no conflict of interests in the present project. The authors also declare that this manuscript has not been previously published, and will not be submitted elsewhere before the decision.
Funding
The research of Yucong Huang was supported in part by the UK Engineering and Physical Sciences Research Council Award EP/L015811/1. The research of Shinya Nishibata was supported in part by Grant-in-Aid for Scientific Research (C)(2) 14540200 of the Ministry of Education, Culture, Sports, Science and Technology Science and Technology.
Declaration of generative AI in scientific writing
The authors declare that no generative AI is used in the writing of this paper.
Appendix A Appendix
A.1 Hardy’s inequality